Find the equation of the radical axis of the two circles x2 + y2 - 4x - 6y - 3 = 0 and x2 + y2 - 12x - 14y + 65 = 0.
a. 2x - 2y + 17 = 0 ​
​b. 2x + 2y + 17 = 0
​ ​c. 2x - 2y - 17 = 0 ​ ​ ​
​d. 2x + 2y - 17 = 0

Answer by greenestamps(13200)



NB: Use "^" (shift-6) to denote exponentiation -- e.g., "x^2" instead of "x2".
Subtract the second equation from the first:

   x^2 + y^2 -  4x -  6y -  3 = 0
   x^2 + y^2 - 12x - 14y + 65 = 0
  -------------------------------
                8x +  8y - 68 = 0

Simplify (divide by 4)

  2x + 2y - 17 = 0

ANSWER d.